Solve by using the quadratic formula.
5x^2 + x = 3

Answer by Edwin McCravy(20054)

Solve by using the quadratic formula.
5x² + x = 3



              5x² + x = 3

Get 0 on the right by subtracting 3 from both sides


           5x² + x - 3 = 0

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where a = 5; b = 1; c = -3

                     _____________
             -(1) ± Ö(1)²-4(5)(-3)
        x = ————————————————————————
                     2(5) 
                   ____ 
             -1 ± Ö1+60
        x = —————————————
                 10

                   __ 
             -1 ± Ö61
        x = ———————————
                10 
                      
Using the + and a calculator, one solution equals 
about .6810249878.
                       
Using the - and a calculator, the other solution 
equals about -.8810249676.

Edwin